OpenStudy (anonymous):

Simplify : (2ab^2)^2(-3a^2b)^2 . AnyOne Know How Too Simplify This ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

36a^6b^6

OpenStudy (anonymous):

36a^6b^6

OpenStudy (anonymous):

36a^6b^6

OpenStudy (anonymous):

apparently several people do!

OpenStudy (anonymous):

How Did You Do It ? I Need Too Know And So I Can Show My Work .

OpenStudy (anonymous):

(2ab^2)^2 u turn it into 4a^b^2

OpenStudy (anonymous):

(-3a^2b)^2 u turn it into 9a^4b^2

OpenStudy (anonymous):

(4a^b^2 )(9a^4b^2)

OpenStudy (anonymous):

u add exponents of like terms

OpenStudy (anonymous):

36a^(2+4) b^(4+2)

OpenStudy (anonymous):

if u understand thats awesome, if u dont ask ask ask again

OpenStudy (anonymous):

So Multiply The 2 Problems ? The 4a... And 9a.... ?

OpenStudy (anonymous):

yes when multiplying add exponents, when its to the exponent like (a^2)^3=a^6 u multiply exponents bu when multiplying add exponents (a^2 )(a^3)=a^5

OpenStudy (anonymous):

Okaa , I Think I UnderStand But What's A Way Of Remembering When Too Add Or Multiply The Exponents ?

OpenStudy (anonymous):

when its u are multiplying variables add exponents when exponent to the is variable like (a^x)^n multiply x*n

OpenStudy (anonymous):

Can You Show Me Another Example Please ?